OpenStudy (anonymous):

help please? how do i convert Convert 4.50 x 105 ng to Mg

OpenStudy (joannablackwelder):

I would suggest to first convert to grams, then to Mg.

OpenStudy (joannablackwelder):

Any ideas of how to do the first step?

OpenStudy (anonymous):

(1. 450000ng * 10^-9grams 1ng = .0045grams 2.) /0045grams * 1Mg 10^6 grams =4.5*10^-10Mg

OpenStudy (jimdemers):

Assuming thast your original number is \[4.50 \times 10^{5} ng\] and that you want the number of mg, all you need to do is reduce the exponent by 6. (There are 1,000,000 ng in a mg.) That leaves you with \[4.50 \times 10^{-1} mg\].